# Consistent Story Generation: Unlocking the Potential of Zigzag Sampling

### v4

@@ 5 Results
Qualitative Comparison. Figures 4 and 6 present qualitative comparisons of our method against existing approaches. For the SDXL model, we compare our method with IP-Adapter [24], ConsisTory [28], StoryDiffusion [27], and 1Prompt1Story [29]. For the FLUX model, we evaluate our method against the training-free, state-of-the-art 1Prompt1Story approach. As shown in Figure 4, IP-Adapter tends to overemphasize identity consistency at the expense of faithfully representing the input text. Meanwhile, ConsisTory, StoryDiffusion, and 1Prompt1Story struggle to maintain subject consistency across different images. In contrast, our method achieves a well-balanced performance, effectively preserving subject identity while accurately following the text descriptions. Figure 6 further underscores the superiority of our approach over the 1Prompt1Story model in terms of both subject consistency and text alignment. Additionally, it highlights the generalizability of our method, which maintains strong performance across diverse network architectures.
@@
Other Applications. Similar to the One-Prompt-One-Story approach [29], our method can be naturally extended to generate long stories containing an arbitrary number of images. This is achieved by applying a sliding-window strategy over the sequence of prompts, where each window of prompts is processed using our proposed method to ensure coherent subject and style consistency across adjacent images. By iteratively moving the window through the entire prompt sequence, our approach can maintain both local and global narrative consistency, enabling the generation of super-long visual stories. A visualization of such an extended story is provided in the appendix, demonstrating the capability of our method to handle narratives of considerable length while preserving high-quality and consistent image generation.
@@ Appendix D Implementation Details
We implement our method on two open-source models: SDXL and FLUX. For SDXL, we use the stabilityai/stable-diffusion-xl-base-1.0 version, and for FLUX, we adopt the black-forest-labs/FLUX.1-dev version. All baseline methods—including IP-Adapter [24], Consistory [28], StoryDiffusion [27], and 1Prompt1Story [29]—are reproduced using their official open-source implementations with default hyperparameters. Since IP-Adapter is designed for image-conditional generation, we adapt it to our setting by first generating an identity image using SDXL with the given identity prompt. This generated image is then used as the conditioning input for IP-Adapter to produce images guided by different prompts.
@@ Appendix E More Details about User Study
We conducted a user study to evaluate our method in comparison with four existing approaches: IP-Adapter [24], Consistory Model [28], Story Diffusion [27], and 1Prompt1Story [29]. All models were used to generate images based on prompts from the ConsiStory+ benchmark, using the same random seeds as reported in their respective papers to ensure a fair comparison.

### v5

@@ 5 Results
Qualitative Comparison. Figures 4 and 6 present qualitative comparisons of our method against existing approaches. For the SDXL model, we compare our method with IP-Adapter [ip-adapter], ConsisTory [consistent-story], StoryDiffusion [storydiffusion], and 1Prompt1Story [one-prompt]. For the FLUX model, we evaluate our method against the training-free, state-of-the-art 1Prompt1Story approach. As shown in Figure 4, IP-Adapter tends to overemphasize identity consistency at the expense of faithfully representing the input text. Meanwhile, ConsisTory, StoryDiffusion, and 1Prompt1Story struggle to maintain subject consistency across different images. In contrast, our method achieves a well-balanced performance, effectively preserving subject identity while accurately following the text descriptions. Figure 6 further underscores the superiority of our approach over the 1Prompt1Story model in terms of both subject consistency and text alignment. Additionally, it highlights the generalizability of our method, which maintains strong performance across diverse network architectures.
@@
Other Applications. Similar to the One-Prompt-One-Story approach [one-prompt], our method can be naturally extended to generate long stories containing an arbitrary number of images. This is achieved by applying a sliding-window strategy over the sequence of prompts, where each window of prompts is processed using our proposed method to ensure coherent subject and style consistency across adjacent images. By iteratively moving the window through the entire prompt sequence, our approach can maintain both local and global narrative consistency, enabling the generation of super-long visual stories. A visualization of such an extended story is provided in the appendix, demonstrating the capability of our method to handle narratives of considerable length while preserving high-quality and consistent image generation.
@@ Appendix D Implementation Details
We implement our method on two open-source models: SDXL and FLUX. For SDXL, we use the stabilityai/stable-diffusion-xl-base-1.0 version, and for FLUX, we adopt the black-forest-labs/FLUX.1-dev version. All baseline methods—including IP-Adapter [ip-adapter], Consistory [consistent-story], StoryDiffusion [storydiffusion], and 1Prompt1Story [one-prompt]—are reproduced using their official open-source implementations with default hyperparameters. Since IP-Adapter is designed for image-conditional generation, we adapt it to our setting by first generating an identity image using SDXL with the given identity prompt. This generated image is then used as the conditioning input for IP-Adapter to produce images guided by different prompts.
@@ Appendix E More Details about User Study
We conducted a user study to evaluate our method in comparison with four existing approaches: IP-Adapter [ip-adapter], Consistory Model [consistent-story], Story Diffusion [storydiffusion], and 1Prompt1Story [one-prompt]. All models were used to generate images based on prompts from the ConsiStory+ benchmark, using the same random seeds as reported in their respective papers to ensure a fair comparison.